# Multi-Armed Bandits with Abstention

### 23 Feb 2024 v1

@@ 3.1 Upper Bounds
Theorem 3.3 below provides two distinct types of theoretical guarantees pertaining to our algorithm’s performance on the cumulative regret ${R^{\mathrm{RG}}_{\mu,c}(T)}$, which is defined in Equation (1) for the fixed-regret setting.
Theorem 3.3.
@@
The comprehensive proof of Theorem 3.3 is deferred to Appendix C.1. The theoretical challenges arising from proving Theorem 3.3 revolve around the quantification of the regret that results from inaccurately estimating the suboptimality gaps associated with the abstention criteria. More precisely, it is crucial to establish upper bounds on $\operatorname{\mathbb{E}}[N_{i}^{(1)}(T)]$ for arms $i$ with $\Delta_{i}<c$ (which, by definition, includes the best arm), and on $\operatorname{\mathbb{E}}[N_{i}^{(0)}(T)]$ for arms $i$ with $\Delta_{i}>c$, from both asymptotic and worst-case perspectives. To accomplish these tasks, we initially decompose the target expectations and subsequently bound the resulting subterms in the two respective regimes. Our approach involves extensive application of analytical techniques inspired by both Thompson Sampling and UCB methodologies.
Furthermore, for the asymptotic upper bound, we take a deeper exploration into the randomized arm sampling dynamics inherent to Less-Exploring Thompson Sampling. A pivotal aspect of this exploration is a characterization of a high probability lower bound on the number of pulls of the optimal arm, as detailed in Lemma C.1, inspired by the work of Korda et al. (2013).
Finally, we remark that the numerous intricacies involved in these analyses preclude us from formulating a generalized strategy akin to the forthcoming Algorithm 2 designed for the fixed-reward setting.
@@ 3.2 Lower Bounds
Definition 3.4 ($R^{\mathrm{RG}}$-consistency).
@@
Now we present both asymptotic and minimax lower bounds on the cumulative regret in Theorem 3.5, which is proved in Appendix C.2.
Theorem 3.5.
@@
Comparing the upper bounds on the cumulative regret of our algorithm FRG-TSwA in Theorem 3.3 with the corresponding lower bounds in Theorem 3.5, it is evident that our algorithm exhibits both asymptotic and minimax optimality.
Asymptotic optimality. For any abstention regret $c>0$ and bandit instance $\mu\in\mathcal{U}$, the regret of our algorithm satisfies the following limiting behaviour:
@@
The above asymptotically optimal result yields several intriguing implications. First, the inclusion of the additional fixed-regret abstention option does not obviate the necessity of differentiating between suboptimal arms and the optimal one, and the exploration-exploitation trade-off remains crucial. In fact, to avoid the case in which the cumulative regret grows polynomially, the agent must still asymptotically allocate the same proportion of pulls to each suboptimal arm, as in the canonical model. This assertion is rigorously demonstrated in the proof of the lower bound (refer to Appendix C.2 for details). Nevertheless, the abstention option does indeed reduce the exploration cost for the agent. Specifically, when exploring any suboptimal arm with a suboptimality gap larger than $c$, our algorithm tends towards employing the abstention option to minimize the instantaneous regret. This aspect is formally established in the proof of the asymptotic upper bound (see Appendix C.1).
Minimax optimality. In the context of worst-case guarantees for the cumulative regret, we focus on the dependence on the problem parameters: $c$, $K$ and $T$. Notably, the $\sum_{i>1}{\Delta_{i}}$ term [^3] This term is unavoidable when the abstention regret $c$ is sufficiently high, since every reasonable algorithm has to allocate a fixed number of pulls to each arm. is typically considered as negligible in the literature (Audibert & Bubeck, 2009; Agrawal & Goyal, 2017; Lattimore & Szepesvári, 2020). Therefore, Theorem 3.3 demonstrates that our algorithm attains a worst-case regret of $O(\sqrt{KT}\wedge cT)$, which is minimax optimal in light of Theorem 3.5.
@@ 4 Fixed-Reward Setting
In this section, we investigate the fixed-reward setting. Here, the reward associated with the abstention option remains consistently fixed at $c\in\mathbb{R}$. When exploring a specific arm, the agent has the capability to determine whether selecting the abstention option yields a higher reward (or equivalently, a lower regret) solely based on its own estimated mean reward. However, in the fixed-regret setting, this decision can only be made by taking into account both its own estimated mean reward and the estimated mean reward of the potentially best arm. In this regard, the fixed-reward setting is inherently less complex than the fixed-regret setting. As a result, it becomes possible for us to design a more general strategy Fixed-Reward Algorithm with Abstention (or FRW-ALGwA), whose pseudocode is presented in Algorithm 2. Despite the straightforward nature of our algorithm, we demonstrate its dual attainment of both asymptotic and minimax optimality through an exhaustive theoretical examination in Sections 4.1 and 4.2.
@@
4:     Choose to abstain ($B_{t}=1$) if and only if
$\hat{\mu}_{A_{t}}(t-1)\leq c.$
@@
[…]
@@ 4.1 Upper Bounds
Recall the definition of the cumulative regret ${R^{\mathrm{RW}}_{\mu,c}(T)}$, as presented in Equation (2) for the fixed-reward setting. Theorem 4.1 establishes both the instance-dependent asymptotic and instance-independent minimax upper bounds for Algorithm 2; see Appendix D.1 for the proof.
Theorem 4.1.
@@
Remark 4.2.
It is worth considering the special case where $c\geq\mu_{1}$, where opting for abstention results in a reward even greater than, or equal to, the mean reward of the best arm. For this particular case, as per Theorem 4.1, since $\mu_{1}\vee c-\mu_{i}\vee c=0$ for all $i>1$, our algorithm achieves a regret of $o(\log T)$. This result, in fact, is not surprising. In contrast to the fixed-regret setting where the regret associated with the abstention option is strictly positive, in this specific scenario of the fixed-reward setting, selecting the abstention option is indeed the optimal action at a single time step, regardless of the arm pulled. Therefore, there is no necessity to distinguish between suboptimal arms and the optimal one, and the exploration-exploitation trade-off becomes inconsequential. However, when the abstention reward is below the mean reward of the best arm, i.e., $c<\mu_{1}$, maintaining a subpolynomial cumulative regret still hinges on the delicate balance between exploration and exploitation, as evidenced by the forthcoming exposition of the asymptotic lower bound.
@@ 4.2 Lower Bounds
We hereby introduce the concept of $R^{\mathrm{RW}}$-consistency for the fixed-reward setting, in a manner analogous to the fixed-regret setting. Following this, we present two distinct lower bounds for the problem of regret minimization in multi-armed bandits with fixed-reward abstention in Theorem 4.4. The proof for Theorem 4.4 is postponed to Appendix D.2.
Definition 4.3 ($R^{\mathrm{RW}}$-consistency).
@@
By comparing the upper bounds in Theorem 4.1 with the lower bounds in Theorem 4.4, it is firmly confirmed that Algorithm 2 is both asymptotically and minimax optimal in the fixed-reward setting.
Asymptotic optimality. For any abstention reward $c\in\mathbb{R}$ and bandit instance $\mu\in\mathcal{U}$, our algorithm ensures the following optimal asymptotic behavior for the cumulative regret:
@@
Since it holds generally that $\mu_{1}\vee c-\mu_{i}\vee c\leq\Delta_{i}$ for each arm $i>1$, our algorithm effectively reduces the cumulative regret in the asymptotic regime through the incorporation of the fixed-reward abstention option.
Minimax optimality. As for the worst-case performance of our algorithm, disregarding the additive term $\sum_{i\in[K]}\left(\mu_{1}\vee c-\mu_{i}\right)$, it achieves an optimal worst-case regret of $O(\sqrt{KT})$. While this worst-case regret aligns with that in the canonical multi-armed bandit model, it is noteworthy that this achievement is non-trivial, demanding meticulous management of the asymptotic regret performance in parallel.
Moreover, there is no occurrence of the phase transition phenomenon in the fixed-reward setting. This absence can be attributed to the intrinsic nature of the fixed-reward abstention option. For any abstention reward $c\in\mathbb{R}$ and online algorithm, we can always construct a challenging bandit instance that leads to a cumulative regret of $\Omega(\sqrt{KT})$, as demonstrated in the proof of the minimax lower bound in Appendix D.2.
@@ 5 Numerical Experiments
In this section, we conduct numerical experiments to empirically validate our theoretical insights. Due to space limitations, we report our results only for the fixed-regret setting here. Results pertaining to the fixed-reward setting are available in Appendix E, where two particular realizations of our algorithm FRW-ALGwA (as outlined in Algorithm 2) are considered. In each experiment, the reported cumulative regrets are averaged over $2,000$ independent trials and the corresponding standard deviations are displayed as error bars in the figures.
Refer to caption (a) Instance $\mu^{\dagger}$ Refer to caption (b) Instance $\mu^{\ddagger}$
@@
To confirm the benefits of incorporating the abstention option, we compare the performance of our proposed algorithm FRG-TSwA (Algorithm 1) with that of Less-Exploring Thompson Sampling (Jin et al., 2023), which serves as a baseline algorithm without the abstention option. We consider two synthetic bandit instances. The first instance $\mu^{\dagger}$ with $K=7$ has uniform suboptimality gaps: $\mu^{\dagger}_{1}=1$ and $\mu^{\dagger}_{i}=0.7$ for all $i\in[K]\setminus\{1\}$. For the second instance $\mu^{\ddagger}$ with $K=10$, the suboptimality gaps are more diverse: $\mu^{\ddagger}_{1}=1$, $\mu^{\ddagger}_{i}=0.7$ for $i\in\{2,3,4\}$, $\mu^{\ddagger}_{i}=0.5$ for $i\in\{5,6,7\}$ and $\mu^{\ddagger}_{i}=0.3$ for $i\in\{8,9,10\}$. The empirical averaged cumulative regrets of both methods with abstention regret $c=0.1$ for different time horizons $T$ are presented in Figure 2. To demonstrate their asymptotic behavior, we also plot the instance-dependent asymptotic lower bound on the cumulative regret (refer to Theorem 3.5) in each sub-figure. It can be observed that FRG-TSwA is clearly superior compared to the non-abstaining baseline, especially for large values of $T$. This demonstrates the advantage of the abstention mechanism. With regard to the growth trend, as the time horizon $T$ increases, the curve corresponding to FRG-TSwA closely approximates that of the asymptotic lower bound. This suggests that the expected cumulative regret of FRG-TSwA matches the lower bound asymptotically, thereby substantiating the theoretical results presented in Section 3.
Refer to caption (a) Instance $\mu^{\dagger}$ Refer to caption (b) Instance $\mu^{\ddagger}$
@@ 6 Conclusions and Future Work
In this paper, we consider, for the first time, a multi-armed bandit model that allows for the possibility of abstaining from accepting the stochastic rewards, alongside the conventional arm selection. This innovative framework is motivated by real-world scenarios where decision-makers may wish to hedge against highly uncertain or risky actions, as exemplified in clinical trials. Within this enriched paradigm, we address both the fixed-regret and fixed-reward settings, providing tight upper and lower bounds on asymptotic and minimax regrets for each scenario. For the fixed-regret setting, we thoughtfully adapt a recently developed asymptotically and minimax optimal algorithm by Jin et al. (2023) to accommodate the abstention option while preserving its attractive optimality characteristics. For the fixed-reward setting, we convert any asymptotically and minimax optimal algorithm for the canonical model into one that retains these optimality properties when the abstention option is present. Finally, experiments on synthetic datasets validate our theoretical results and clearly demonstrate the advantage of incorporating the abstention option.
As highlighted in Remark 3.6, a fruitful avenue for future research lies in expanding the abstention model from $K$-armed bandits to linear bandits. An intriguing inquiry is whether the inclusion of the abstention feature can lead to enhanced asymptotic and minimax theoretical guarantees. Furthermore, in our work, the abstention option exerts no influence on the stochastic observation from the selected arm. Delving into more sophisticated and general approaches to model the effect of the abstention option promises as a captivating future direction.

### 22 Mar 2026 v2

@@ 3.1 Upper Bounds
Theorem 1 below provides two distinct types of theoretical guarantees pertaining to our algorithm’s performance on the cumulative regret $R^{\mathrm{RG}}_{\mu,c}(T)$}, defined in Equation (1) for the fixed-regret setting.
Theorem 1.
@@
The proof of Theorem 1 is deferred to Appendix D.1. The theoretical challenges arising from proving Theorem 1 revolve around the quantification of the regret that results from inaccurately estimating the suboptimality gaps associated with the abstention criteria. More precisely, it is crucial to establish upper bounds on $\operatorname{\mathbb{E}}[N_{i}^{(1)}(T)]$ for arms $i$ with $\Delta_{i}<c$ (which, by definition, includes the best arm), and on $\operatorname{\mathbb{E}}[N_{i}^{(0)}(T)]$ for arms $i$ with $\Delta_{i}>c$. These bounds need to be derived from both asymptotic and minimax perspectives, adding layers of complexity to the analytical process. To accomplish these tasks, we initially decompose the target expectations into manageable components, and subsequently bound the resulting subterms in the two respective regimes. Our approach, merging TS-based sampling rule with LCB-based abstention criteria, necessitates a careful amalgamation of both TS-type and UCB-type analytical techniques.
Furthermore, for the asymptotic upper bound, we take a deeper exploration into the randomized arm sampling dynamics inherent to Less-Exploring Thompson Sampling. A pivotal aspect of this exploration is a characterization of a high probability lower bound on the number of pulls of the optimal arm, as detailed in Lemma 5, inspired by the work of Korda et al. (2013).
Finally, we remark that the numerous intricacies involved in these analyses preclude us from formulating a generalized strategy akin to the forthcoming Algorithm 2 for the fixed-reward setting.
@@ 3.2 Lower Bounds
Definition 1 ($R^{\mathrm{RG}}$-consistency).
@@
Now we present both asymptotic and minimax lower bounds on the cumulative regret in Theorem 2, which is proved in Appendix D.2.
Theorem 2.
@@
Comparing the upper bounds on the cumulative regret of our algorithm FRG-TSwA in Theorem 1 with the corresponding lower bounds in Theorem 2, it is evident that our algorithm exhibits both asymptotic and minimax optimality.
Asymptotic optimality.
For any abstention regret $c>0$ and bandit instance $\mu\in\mathcal{U}$, the regret of our algorithm satisfies the following limiting behaviour:
@@
The above asymptotically optimal result yields several intriguing implications. First, the inclusion of the additional fixed-regret abstention option does not obviate the necessity of differentiating between suboptimal arms and the optimal one, and the exploration-exploitation trade-off remains crucial. In fact, to avoid the case in which the cumulative regret grows polynomially, the agent must still asymptotically allocate the same proportion of pulls to each suboptimal arm, as in the canonical model. This assertion is rigorously demonstrated in the proof of the lower bound (refer to Appendix D.2 for details). Nevertheless, the abstention option does indeed reduce the exploration cost for the agent. Specifically, when exploring any suboptimal arm with a suboptimality gap larger than $c$, our algorithm tends towards employing the abstention option to minimize the instantaneous regret. This aspect is formally established in the proof of the asymptotic upper bound (see Appendix D.1).
Minimax optimality.
In the context of worst-case guarantees for the cumulative regret, we focus on the dependence on $c$, $K$, and $T$. Notably, the $\sum_{i>1}{\Delta_{i}}$ term [^4] This term is unavoidable when the abstention regret $c$ is sufficiently high, since every reasonable algorithm has to allocate a fixed number of pulls to each arm. is typically considered as negligible in the literature (Audibert & Bubeck, 2009; Agrawal & Goyal, 2017; Lattimore & Szepesvári, 2020). Therefore, Theorem 1 demonstrates that our algorithm attains a worst-case regret of $O(\sqrt{KT}\wedge cT)$, which is minimax optimal in light of Theorem 2.
@@
Stochastic and heterogeneous abstention regret.
If the regret incurred by the agent when opting for abstention is not deterministic but stochastically generated from a distribution with known expectation $c$, Algorithm 1 remains effective, and our analyses of the upper and lower bounds remain valid. Additionally, another natural extension of our model is that the abstention regret is different for different arms. Specifically, the agent incurs a regret of $c_{i}>0$ when pulling arm $i$ and choosing abstention. In this scenario, the abstention criteria in Algorithm 1 can be adapted accordingly, and the same asymptotic and minimax optimality can be established; see Appendix D.3 for details.
@@ 4 Fixed-Reward Setting
In this section, we investigate the fixed-reward setting. Here, the reward associated with the abstention option remains consistently fixed at $c\in\mathbb{R}$. When exploring a specific arm, the agent has the capability to determine whether selecting the abstention option yields a higher reward solely based on its own estimated mean reward. However, in the fixed-regret setting, this decision can only be made by taking into account both its own estimated mean reward and the estimated mean reward of the potentially best arm. In this regard, the fixed-reward setting is inherently less complex than the fixed-regret setting. As a result, it becomes possible for us to design a more general strategy Fixed-Reward Algorithm with Abstention (or FRW-ALGwA), whose pseudocode is presented in Algorithm 2. Despite the straightforward nature of our algorithm, we demonstrate its dual attainment of both asymptotic and minimax optimality through an exhaustive theoretical examination in Sections 4.1 and 4.2.
@@
4:  Choose to abstain ($B_{t}=1$) if and only if $\hat{\mu}_{A_{t}}(t-1)\leq c.$
@@
FRW-ALGwA leverages a base algorithm ALG that is asymptotically and minimax optimal for canonical multi-armed bandits as its input. For comprehensive definitions of asymptotic and minimax optimality within the canonical model, see Appendix A. Notably, eligible candidate algorithms include $\text{KL-UCB}^{++}$ (Ménard & Garivier, 2017), ADA-UCB (Lattimore, 2018), MOTS-$\mathcal{J}$ (Jin et al., 2021) and Less-Exploring Thompson Sampling (Jin et al., 2023). In our algorithm, at each time step $t$, the base algorithm determines the selected arm $A_{t}$ according to the partial interaction historical information $(A_{1},X_{1},A_{2},X_{2},\ldots,A_{t-1},X_{t-1})$. Subsequently, the algorithm decides whether or not to abstain, indicated by the binary random variable $B_{t}$, by comparing the empirical mean of the arm $A_{t}$, denoted as $\hat{\mu}_{A_{t}}(t-1)$, to the abstention reward $c$.
@@ 4.1 Upper Bounds
Recall the definition of the cumulative regret $R^{\mathrm{RW}}_{\mu,c}(T)$}, as presented in Equation (2) for the fixed-reward setting. Theorem 3 establishes both the instance-dependent asymptotic and instance-independent minimax upper bounds for Algorithm 2; see Appendix E.1 for the proof.
Theorem 3.
@@
Remark 6.
It is worth considering the special case where $c\geq\mu_{1}$, where opting for abstention results in a reward even greater than, or equal to, the mean reward of the best arm. For this particular case, as per Theorem 3, since $\mu_{1}\vee c-\mu_{i}\vee c=0$ for all $i>1$, our algorithm achieves a regret of $o(\log T)$. This result, in fact, is not surprising. In contrast to the fixed-regret setting where the regret associated with the abstention option is strictly positive, in this specific scenario of the fixed-reward setting, selecting the abstention option is indeed the optimal action at a single time step, regardless of the arm pulled. Therefore, there is no necessity to distinguish between suboptimal arms and the optimal one, and the exploration-exploitation trade-off becomes inconsequential. However, when the abstention reward is below the mean reward of the best arm, i.e., $c<\mu_{1}$, maintaining a subpolynomial cumulative regret still hinges on the delicate balance between exploration and exploitation, as evidenced by the forthcoming exposition of the asymptotic lower bound.
@@ 4.2 Lower Bounds
We hereby introduce the concept of $R^{\mathrm{RW}}$-consistency for the fixed-reward setting, in a manner analogous to the fixed-regret setting. Following this, we present two distinct lower bounds for the problem of regret minimization in multi-armed bandits with fixed-reward abstention in Theorem 4. The proof for Theorem 4 is postponed to Appendix E.2.
Definition 2 ($R^{\mathrm{RW}}$-consistency).
@@
By comparing the upper bounds in Theorem 3 with the lower bounds in Theorem 4, it is firmly confirmed that Algorithm 2 is both asymptotically and minimax optimal in the fixed-reward setting.
Asymptotic optimality.
For any abstention reward $c\in\mathbb{R}$ and bandit instance $\mu\in\mathcal{U}$, our algorithm ensures the following optimal asymptotic behavior for the cumulative regret:
@@
Since it holds that $\mu_{1}\vee c-\mu_{i}\vee c\leq\Delta_{i}$ for all $i>1$, our algorithm effectively reduces the cumulative regret in the asymptotic regime through the incorporation of the fixed-reward abstention option.
Minimax optimality.
Concerning the worst-case performance of our algorithm, disregarding the additive term $\sum_{i\in[K]}\left(\mu_{1}\vee c-\mu_{i}\right)$, it achieves an optimal worst-case regret of $O(\sqrt{KT})$. While this worst-case regret is the same as that for canonical multi-armed bandits, achieving it is non-trivial as we have to simultaneously achieve the minimal asymptotic instance-dependent regret.
Moreover, there is no occurrence of the phase transition phenomenon in the fixed-reward setting. This absence can be attributed to the intrinsic nature of the fixed-reward abstention option. For any abstention reward $c\in\mathbb{R}$ and online algorithm, we can always construct a challenging bandit instance that leads to a cumulative regret of $\Omega(\sqrt{KT})$, as demonstrated in the proof of the minimax lower bound in Appendix E.2.
@@ 5 Numerical Experiments
In this section, we conduct numerical experiments to empirically substantiate our theoretical insights. To reduce clutter, we report our results only for the fixed-regret setting here. Results pertaining to the fixed-reward setting are deferred to Appendix F, where we consider two particular realizations of our algorithm based on Less-Exploring Thompson Sampling (Jin et al., 2023) and $\text{KL-UCB}^{++}$ (Ménard & Garivier, 2017). In each experiment, the reported cumulative regrets are averaged over $2,000$ independent trials and the corresponding standard deviations are displayed as error bars in the figures.
To confirm the benefits of incorporating the abstention option, we compare the performance of our proposed algorithm FRG-TSwA (Algorithm 1) with that of Less-Exploring Thompson Sampling (Jin et al., 2023), which serves as a baseline algorithm without the abstention option. We consider two synthetic bandit instances. The first instance $\mu^{\dagger}$ with $K=7$ has uniform suboptimality gaps: $\mu^{\dagger}_{1}=1$ and $\mu^{\dagger}_{i}=0.7$ for all $i\in[K]\setminus\{1\}$. For the second instance $\mu^{\ddagger}$ with $K=10$, the suboptimality gaps are more diverse: $\mu^{\ddagger}_{1}=1$, $\mu^{\ddagger}_{i}=0.7$ for $i\in\{2,3,4\}$, $\mu^{\ddagger}_{i}=0.5$ for $i\in\{5,6,7\}$ and $\mu^{\ddagger}_{i}=0.3$ for $i\in\{8,9,10\}$. The empirical averaged cumulative regrets of both methods with abstention regret $c=0.1$ for different time horizons $T$ are presented in Figure 2. To demonstrate their asymptotic behavior, we also plot the instance-dependent asymptotic lower bound on the cumulative regret (refer to Theorem 2) in each sub-figure. It can be observed that FRG-TSwA is clearly superior compared to the non-abstaining baseline, especially for large values of $T$. This demonstrates the advantage of the abstention mechanism. With regard to the growth trend, as the time horizon $T$ increases, the curve corresponding to FRG-TSwA closely approximates that of the asymptotic lower bound. This suggests that the expected cumulative regret of FRG-TSwA matches the lower bound asymptotically, thereby substantiating the theoretical results presented in Section 3.
(a) Instance $\mu^{\dagger}$
(b) Instance $\mu^{\ddagger}$
@@
(a) Instance $\mu^{\dagger}$
(b) Instance $\mu^{\ddagger}$
@@ 6 Conclusions and Future Work
In this paper, we consider, for the first time, a multi-armed bandit model that allows for the possibility of abstaining from accepting the stochastic rewards, alongside the conventional arm selection. This innovative framework is motivated by real-world scenarios where decision-makers may wish to hedge against highly uncertain or risky actions, as exemplified in clinical trials. Within this enriched paradigm, we address both the fixed-regret and fixed-reward settings, providing tight upper and lower bounds on asymptotic and minimax regrets for each scenario. For the fixed-regret setting, we thoughtfully adapt a recently developed asymptotically and minimax optimal algorithm by Jin et al. (2023) to accommodate the abstention option while preserving its attractive optimality characteristics. For the fixed-reward setting, we convert any asymptotically and minimax optimal algorithm for the canonical model into one that retains these optimality properties when the abstention option is present. We remark that our proposed algorithms are computationally efficient, incurring only a negligible constant-time overhead per step beyond the standard arm selection mechanisms. Finally, experiments on synthetic datasets validate our theoretical results and clearly demonstrate the advantage of incorporating the abstention option.
While our analysis focuses on Gaussian rewards, the framework naturally extends to other distributions, provided that specific boundary cases are handled with care. As highlighted in Remark 5, a fruitful avenue for future research lies in expanding the abstention model from $K$-armed bandits to linear bandits; specifically, it remains an open question whether the inclusion of the abstention feature can lead to enhanced asymptotic and minimax theoretical guarantees. Furthermore, in our work, the abstention option exerts no influence on the stochastic observation from the selected arm. Delving into more sophisticated and general approaches to model the effect of the abstention option promises as a captivating future direction.
Acknowledgments
T. Jin and V. Y. F. Tan are supported by a Singapore Ministry of Education (MOE) AcRF Tier 2 grant under grant number A-8004062-00-00.